# How to Build a Quantum Supercomputer: Scaling Challenges and Opportunities

### v1

@@ A.3 Propagation of errors in qubitization
In the qubitization approach, there are three main sources of error, as in von Burg et al. (2021a). The first arises from truncating the small eigenvalues of the double-factorized Hamiltonian, but by keeping track of the value of the truncated eigenvalues it is possible to bound the 2-norm difference between the original Hamiltonian an the truncated Hamiltonian. The second source of error arises from approximations in the implementation of the qubitization operator, and can be decomposed into two terms. Namely, when approximating coefficients used during the summing of operators in the LCU decompositions, as well as in the rotation angles in the diagonalization operations in the innermost decomposition of the double factorized Hamiltonian with finite bits of precision. However, by increasing the number of bits used to implement these operations the error terms can be driven to zero. The final contribution to the error is the imprecision in the QPE, as in the PF approach, but this error and requisite number of logical ancillary qubits can be bounded using the total number of repetitions of the qubitization operator.
@@
At the logical level, it is possible to breakdown the error in QPE of the qubitization approach as in Equation (23) of Babbush et al. (2018). Namely, the output energy of the QPE is within
@@
$\Delta E\leq\lambda\sqrt{\left(\frac{\pi}{2^{m}}\right)^{2}+(\epsilon_{H}+\pi\epsilon_{\text{QPE}})^{2}}$ (34)
of the Hamiltonian used in the qubization approach, where $\lambda$ is the 1-norm of the Hamiltonian, $m$ is the number of bits in the QPE, $\epsilon_{\text{QPE}}$ is related to the error in implementing the QPE and is usually negligable, and $\epsilon_{H}$ is related to the 2-norm difference between the qubitization operators of the double factorized Hamiltonian and its implementation.
Following the analysis from Babbush et al. (2018), we can bound the error contributed from the finite bit precision approximations. The error in the qubitization operator is given by
$\epsilon$ $\leq\|e^{i\arccos(H/\lambda)}-e^{i\arccos(\tilde{H}/\lambda)}\|$
$\leq\|\arccos(H/\lambda)-\arccos(\tilde{H}/\lambda)\|$
$\leq\sum_{p=0}^{\infty}\frac{(2p-1)!!}{\lambda^{2p+1}(2p+1)(2p)!!}\|H^{2p+1}-\tilde{H}^{2p+1}\|$ (35)
The term $\|H^{2p+1}-\tilde{H}^{2p+1}\|$ can be bounded by $(2p+1)(\|H-\tilde{H}\|+\|H\|)^{2p}(\|H-\tilde{H}\|)$, giving us
$\epsilon$ $\leq\sum_{p=0}^{\infty}\frac{(2p-1)!!}{\lambda^{2p+1}(2p+1)(2p)!!}(2p+1)(\|H-\tilde{H}\|)^{2p+1}$
$\leq\frac{\|H-\tilde{H}\|}{\lambda}\left[1-\left(\frac{\|H\|+\|H-\tilde{H}\|}{\lambda}\right)^{2}\right]^{1/2}.$ (36)
Let $\Gamma\equiv\|H-\tilde{H}\|$. Solving for $\Gamma$ in the above inequality, we obtain
@@
$\Gamma$ $\leq\frac{\sqrt{2}\Delta E}{4\left(1+\frac{\Delta E^{2}}{8\lambda^{2}}\right)}\left(1-\frac{\|H\|^{2}}{\lambda^{2}}\right),$ (37)
Thus, if we aim the error contribution from implementing qubitization to be within $\Delta E$, then we must choose the bit precisions to be large enough that the 2-norm difference satisfies the above.
Appendix B Quantum resource estimations
@@
Once a target logical quantum circuit has been generated, we construct a fault-tolerant architecture that can implement this circuit to conduct a QRE analysis using the TopQAD toolkit 1QB Information Technologies (2024) (1QBit). The software’s approach to creating a fault-tolerant architecture is described in Ref. Silva et al. (2024a). For a given quantum circuit and success probability, we generate an architecture that would feasibly run the computation at the requisite precision. This architecture is one that allows us to estimate the resources required for a specific quantum circuit using hardware that can implement a rotated surface code layout. By abstracting the hardware away, we are able to focus on the layout and from there construct an architecture that can be used to implement those operations required for FTQC.
@@ B.1 The compilation process
The main idea behind this construction is to transform the given circuit into an optimized sequence of $\pi/8$ Pauli rotations, and then to process these rotations using multi-qubit lattice surgery to connect distant qubits (Litinski, 2019; Beverland et al., 2022; Silva et al., 2024b). These $\pi/8$ Pauli rotations can then be implemented on the underlying architecture, where magic states are distilled and consumed through specific applications of lattice surgery. This procedure results in a very nontrivial simultaneity condition, as the bus qubits used in the lattice surgeries can only be used for a single rotation at a given point in time. Additionally, there are several conditions for how the bus qubits can interact with qubits storing data for the circuit, leading to complications in the underlying architecture. However, once these conditions have been taken into account, various classical scheduling processes can be used to generate the necessary schedule of operations that we then implement via lattice surgery.
@@
Given that the architecture considered requires that quantum operations are represented by Pauli rotations, the next step is to transpile the circuit consisting of Clifford and $T$ gates into a circuit consisting of Pauli rotations. The circuit described in the Clifford + $T$ gate set is first converted to a sequence of $\pi/4$ (Clifford) and $\pi/8$ (non-Clifford) Pauli rotations according to the conversion rules described in Ref. Litinski (2019). After conversion, a procedure is run to remove the Clifford operations from the circuit using commutation rules, leaving only $\pi/8$ rotations. This procedure can be run efficiently using the symplectic representation of Clifford gates Silva et al. (2024b). Additionaly, since commutable $\pi/8$ rotations can be reordered such that adjacent $\pi/8$ rotations with same axis of rotation can be combined, this allows us to perform operations corresponding to multiple rotation commutations in a single step effectively reducing the $T$ count. As discussed in Ref. Silva et al. (2024b), this transpilation procedure drastically decreases the overall running time of the circuit even if the resulting circuit makes the operations less parallelizable due to its increased density.
@@ B.2 The assembly process
At this point, we have constructed a logical circuit tailored to an implementation on a surface code encoding logical qubits. The next step in the pipeline to generate the QREs is the assemble of the structures required for FTQC that will allow the scheduling of the $\pi/8$ Pauli rotations in the circuit. As illustrated in fig. 14, the architecture considered for the scheduling of the logical operations features a core processor, comprising a memory fabric with two-tile two-qubit patches of data qubits and an auto-correcting buffer, which is connected to the MSF using bus qubits, mirroring the configuration used in Ref. Litinski (2019).
@@
Central to our approach is the utilization of a multi-level MSF for magic state distillation, where the fidelity of magic states undergoes iterative enhancement across successive distillation levels. Low-fidelity magic states are created from operations on physical qubits at magic state preparation units following a magic state preparation protocol Gidney (2023); Gidney et al. (2024) as described in III.5. Then, at each distillation level, the MSF used the lower-fidelity magic states to create higher-fidelity magic states that are dispatched to a dedicated area where magic states can be enlarged to the required code distance that interfaces with the next round of distillation. A 15:1 distillation protocol is assumed to be used by the distillation units at all levels due to its capacity to improve magic state fidelity in $O(P_{T}^{3})$, where $P_{T}$ is the logical error rate for input magic states Beverland et al. (2022). The uppermost level of the MSF connects to the memory fabric via a buffer space that allows magic states to be temporarily stored before being consumed within the memory fabric. This space is designed to incorporate auto-correcting buffers, named for their capability to execute corrective measures concurrently with magic state consumption, notably enabling the auto-correcting of $\pi/8$ operations. The auto-correcting buffers and the memory fabric comprise the core processor of the device.
@@
The scheduling methodology employed in the studied topological architecture is presented in Ref. Silva et al. (2024b) and addresses the sequencing of operations and the allocation of logical resources required for establishing connections between distant qubits in the core processor as needed. Due to the reduced parallelization potential of the $\pi/8$ operations, we assume a serial scheduling is employed in the QREs presented here. If nonrestrictive availability of magic states is ensured, and considering that the expected time to execute a $\pi/8$ rotation is equal to one logical cycle because of the auto-correcting buffers, the minimum number of logical cycles required to execute the entire circuit in a serial scheduling is equal to its $T$ count, ignoring the warm-up time. Each logical cycle requires performing $d_{core}$ parity checks, where $d_{core}$ is the code distance of the logical qubits in the core processor, each taking a time $T_{M}+4T_{2}+2T_{1}+T_{M}+t_{R}$, considering the measurement time $t_{M}$, the reset time $t_{R}$, the single-qubit gate time $t_{1}$, and the two-qubit gate time $t_{2}$. Therefore, while it is easy to generate time estimates for circuits scheduled in serial, generating the space estimates require creating a MSF with enough distillation units capable of distilling magic states quickly enough to keep the core processor constantly busy.
@@
The errors of the core processor and the MSF are modeled and predicted following the pipeline described in Silva et al. (2024a). Since we provide QREs for the case with a never idling core processor, the accumulated errors in the core are only resulting from the Clifford operations required for the multi-qubit lattice surgeries performed and the protection of the idling data qubits while lattice surgeries involving other data qubits are occurring in the core. The accumulated errors in the core processor is approximated as
$E_{\text{core}}\approx(2Q+\sqrt{8Q}+29)Te_{\text{mem},\text{core}}.$ (39)
which assumes that all $(2Q+\sqrt{8Q}+29)$ logical qubits in the core processor (approximated size of the memory fabric and buffer) following the design presented in Figure 14, where $Q$ is the number of data qubits in the circuit, are susceptible to result in an error with probability $e_{\text{mem},\text{core}}$ during all the $T$ logical cycles required to run the circuit. The error rate $e_{\text{mem},\text{core}}$ is derived from emulations of the FTQC protocol for quantum memory. These emulations establish the correlation between code distance and logical error rates based on a given choice of physical parameters, resulting in a predictive model obtained by regression from numerical simulations at low code distances, using efficient stabilizer circuit simulators Gidney (2021) following the descriptions in Section III.2. In the MSF, the error rate of output magic states is resulting from the preparation, distillation and expansion procedures. Following Silva et al. (2024a), considering $e_{\text{prep}}$ as the error rate for the magic states prepared from physical qubits and that the Clifford and growth accumulated errors can be approximated to the memory errors, i.e., $e_{\text{cliff}}=e_{\text{grow}}=e_{\text{mem}}$, the magic state error rates of the entire MSF using 15:1 distillation units can be calculated recursively as follows:
@@
Therefore, the error rate of the magic state input to the core processor is $e_{\text{core}}=e_{\text{in},L+1}$ for an MSF with $L$ distillation levels. Given that $T$ magic states needs to be distilled for the entire execution of the quantum program, we have that
@@
The choice of hardware parameters to determine the required number of distillation levels $L$ and code distances $d_{l},\forall l\in{1,\ldots,L+1}$, which includes the core processor as $l=L+1$, is such that it must reach the target logical error rates based on Equations 38 to 40. The process followed to make these decisions is described in Ref. Silva et al. (2024a). In summary, the core processor’s code distance $d_{L+1}$ is minimized, assuming $E_{\text{msf}}=0$. Then, it sets the first level code distance $d_{1}$ considering the residual error budget left after the core level logical encoding is decided, i.e., $E_{\text{msf}}\leq E-E_{\text{core}}$. Next, it determines the number of distillation levels $L$ required to meet the magic state error rate requirement $e_{core}$ derived from 40 for the residual error budget. Finally, if $L>1$, it calculates the code distances $d_{l}$ for all levels that can meet the error budget using the minimum number of physical qubits across the whole architecture. Once these decisions have been made, the assembler determines the number of distillation units required for a steady flow of magic states to the core processor such that the distillation rate of magic states output from the MSF matches the consumption rate of magic states in the core processor.

### v2

@@ A.3 Propagation of errors in qubitization
In this section, we discuss the various errors incurred in the process of generating the logical quantum circuits for the double-factorized (DF) qubitization algorithm, and how we can guarantee electronic-structure quantum computations to estimate the Hamiltonian eigenvalues to a given target accuracy by satisfying certain error bounds within predetermined error budgets. In QPE based on qubitization, the spectrum of the molecular Hamiltonian $H$ is sampled by measuring the phase that is accumulated on an initial input quantum state through multiple controlled executions of the quantum walk operator $e^{i\arccos(H/\lambda)}$ (or $e^{i\arcsin(H/\lambda)}$) in place of the time-evolution operator $\exp(-iHt)$, see Refs. Berry et al. (2018); Poulin et al. (2018). In the DF qubitization algorithm, there are four main sources of error (see Refs. von Burg et al. (2021a); Lee et al. (2021)):
  1. 1.
    Error associated with the truncation in the double-factorization procedure;
  2. 2.
    Error associated with the binary approximation of the Hamiltonian coefficients when implementing the prepare oracle;
  3. 3.
    Error associated with the approximation of the individual Givens rotations in implementing the basis rotations; and
  4. 4.
    Error associated with the measurement readout in phase estimation.
@@
The first source of error arises from truncating the small eigenvalues of the double-factorized Hamiltonian, but by keeping track of the values of the truncated eigenvalues it is possible to bound the 2-norm of the difference between the original Hamiltonian and the truncated Hamiltonian. The second and third sources of error arise from approximations in the implementation of the LCU oracles associated with the qubitization operator, namely, binary approximations of the coefficients involved in the prepare operation as well as binary approximations of the rotation angles in the diagonalization operations involved in the innermost decomposition of the double-factorized Hamiltonian (with finite bits of precision). The final contribution to the error is the imprecision in the measurement readout of QPE; this error and the requisite number of logical ancillary qubits can be bounded using the total number of repetitions of the qubitization operator. For chemical significance, the total overall target error in estimating the ground-state energy should be at most 1.6 millihartrees, that is, the overall error budget for estimating the lowest eigenvalue of the molecular Hamiltonian is $\epsilon=1.6$ mHa (or the much lower target error $\epsilon=0.1\text{ mHa}$ for quantitative accuracy).
@@
At the logical level, it is possible to break down the error in QPE of the qubitization approach as in Equation (23) of Ref. Babbush et al. (2018), namely, the output energy of the QPE is within
@@
$\Delta E\leq\lambda\sqrt{\left(\frac{\pi}{2^{m}}\right)^{2}+(\epsilon_{H}+\pi\epsilon_{\text{\tiny QFT}})^{2}}$ (36)
of the Hamiltonian used in the qubitization approach. Here, $\lambda$ is the 1-norm of the Hamiltonian (computed as the sum of the absolute values of the Hamiltonian weightings), $m$ is the number of bits in the QPE, $\epsilon_{\text{\tiny QFT}}$ is related to the error in implementing the quantum Fourier transform and is usually negligable, and $\epsilon_{H}$ is the systematic error in the phase resulting from the gate synthesis in implementing the qubitization operator of the double-factorized Hamiltonian. According to Equation 36, if we aim to estimate spectra to within error $\Delta E$, we can achieve this by choosing (see Babbush et al. (2018)):
@@
$m=\left\lceil\log_{2}\left(\frac{\pi\lambda}{\sqrt{2}\Delta E}\right)\right\rceil,\;\epsilon_{H}\leq\frac{\sqrt{2}\Delta E}{4\lambda},\;\epsilon_{\text{\tiny QFT}}\leq\frac{\sqrt{2}\Delta E}{4\pi\lambda}.$ (37)
Following the analysis outlined in the Appendix of Ref. Babbush et al. (2018), we can bound the error contributions arising from finite-bit precision approximations of the quantum circuit encoding of the double-factorized Hamiltonian. Let $H=\sum_{\ell=0}^{L-1}w_{\ell}H_{\ell}$ denote the double-factorized Hamiltonian (after truncation of the terms pertaining to small eigenvalues), and let $\widetilde{H}=\sum_{\ell=0}^{L-1}\widetilde{w}_{\ell}\widetilde{H}_{\ell}$ denote the corresponding approximate encoding of $H$ resulting from finite-bit precision approximations of Hamiltonian coefficients and Givens rotations. Then, the associated propagated error into the eigenphase obeys:
$\epsilon_{H}$ $\leq\|e^{i\arccos(H/\lambda)}-e^{i\arccos(\widetilde{H}/\lambda)}\|$
$\leq\|\arccos(H/\lambda)-\arccos(\widetilde{H}/\lambda)\|$
$\leq\sum_{p=0}^{\infty}\frac{(2p-1)!!}{\lambda^{2p+1}(2p+1)(2p)!!}\|H^{2p+1}-\widetilde{H}^{2p+1}\|.$ (38)
Let $\chi:=\|H-\widetilde{H}\|$. Then the term $\|H^{2p+1}-\widetilde{H}^{2p+1}\|$ can be bounded from above by $(2p+1)(\|H\|+\chi)^{2p}\chi$, which yields
$\epsilon_{H}$ $\leq\sum_{p=0}^{\infty}\frac{(2p-1)!!}{\lambda^{2p+1}(2p)!!}(\|H\|+\chi)^{2p}\chi$
$\leq\frac{\chi}{\lambda}\left[1-\left(\frac{\|H\|+\chi}{\lambda}\right)^{2}\,\right]^{-1/2}.$ (39)
Thus, in order for the error contributions arising from finite-bit precision approximations in implementing the qubitization quantum circuit to satisfy the bound specified in Equation 37, we must choose the bit precisions to be large enough such that
$\chi$ $\leq\frac{\sqrt{2}\Delta E}{4\left(1+\frac{\Delta E^{2}}{8\lambda^{2}}\right)}\left(1-\|H\|^{2}/\lambda^{2}\right).$ (40)
This inequality is used to determine the number of bits required for the binary approximations of the Hamiltonian coefficients $w_{\ell}$ and the binary approximations of the rotation angles associated with the diagonalizing fermionic basis transformations that yield the double-factorized form.
Appendix B Quantum resource estimation
@@
Once a target logical quantum circuit has been generated, we construct a fault-tolerant architecture that can implement this circuit to conduct a QRE analysis using the TopQAD toolkit 1QB Information Technologies (2024) (1QBit). The software’s approach to creating a fault-tolerant architecture is described in Ref. Silva et al. (2024a). For a given quantum circuit and success probability, we generate an architecture that would feasibly run the computation at the requisite precision. This architecture allows us to estimate the resources required for a specific quantum circuit using hardware that can implement a rotated surface code layout. By abstracting the hardware away, we are able to focus on the layout, and from there construct an architecture that can be used to implement those operations required for FTQC.
@@ B.1 The compilation process
The main idea behind this construction is to transform the given circuit into an optimized sequence of $\pi/8$ Pauli rotations, and then to process these rotations using multi-qubit lattice surgery to connect distant qubits (Litinski, 2019; Beverland et al., 2022; Silva et al., 2024b). These $\pi/8$ Pauli rotations can then be implemented on the underlying architecture, where magic states are distilled and consumed through specific applications of lattice surgery. This procedure results in a nontrivial simultaneity condition, as the bus qubits used in the lattice surgeries can only be used for a single rotation at a given point in time. Additionally, there are several conditions for how the bus qubits can interact with qubits storing data for the circuit, leading to complications in the underlying architecture. However, once these conditions have been taken into account, various classical scheduling processes can be used to generate the necessary schedule of operations that we then implement via lattice surgery.
@@
Given that the architecture considered requires that quantum operations are represented by Pauli rotations, the next step is to transpile the circuit consisting of Clifford and $T$ gates into a circuit consisting of Pauli rotations. The circuit described in the Clifford+$T$ gate set is first converted to a sequence of $\pi/4$ (Clifford) and $\pi/8$ (non-Clifford) Pauli rotations according to the conversion rules described in Ref. Litinski (2019). After conversion, a procedure is run to remove the Clifford operations from the circuit using commutation rules, leaving only $\pi/8$ rotations. This procedure can be run efficiently using the symplectic representation of Clifford gates Silva et al. (2024b). Additionally, since commutable $\pi/8$ rotations can be reordered such that adjacent $\pi/8$ rotations with same axis of rotation can be combined, this allows us to perform operations corresponding to multiple rotation commutations in a single step effectively reducing the $T$ count. As discussed in Ref. Silva et al. (2024b), this transpilation procedure drastically decreases the overall running time of the circuit even if the resulting circuit makes the operations less parallelizable due to its increased density.
@@ B.2 The assembly process
At this point, we have constructed a logic circuit tailored to an implementation on a surface code encoding logical qubits. The next step in the pipeline to generate the QREs is the assemble of the structures required for FTQC that will allow the scheduling of the $\pi/8$ Pauli rotations in the circuit. As illustrated in Figure 15, the architecture considered for the scheduling of the logical operations features a core processor, comprising a memory fabric with two-tile two-qubit patches of data qubits and an auto-correcting buffer, which is connected to the MSF using bus qubits, mirroring the configuration used in Ref. Litinski (2019).
@@
Central to our approach is the utilization of a multi-level MSF for magic state distillation, where the fidelity of magic states undergoes iterative enhancement across successive distillation levels. Low-fidelity magic states are created from operations on physical qubits at magic state preparation units following a magic state preparation protocol Gidney (2023); Gidney et al. (2024) as described in Section III.5. Then, at each distillation level, the MSF used the lower-fidelity magic states to create higher-fidelity magic states that are dispatched to a dedicated area where magic states can be enlarged to the required code distance that interfaces with the next round of distillation. A 15:1 distillation protocol is assumed to be used by the distillation units at all levels due to its capacity to improve magic state fidelity in $O(P_{T}^{3})$, where $P_{T}$ is the logical error rate for input magic states Beverland et al. (2022). The uppermost level of the MSF connects to the memory fabric via a buffer space that allows magic states to be temporarily stored before being consumed within the memory fabric. This space is designed to incorporate auto-correcting buffers, named for their capability to execute corrective measures concurrently with magic state consumption, notably enabling the auto-correcting of $\pi/8$ operations. The auto-correcting buffers and the memory fabric comprise the core processor of the device.
@@
The scheduling methodology employed in the studied topological architecture is presented in Ref. Silva et al. (2024b) and addresses the sequencing of operations and the allocation of logical resources required for establishing connections between distant qubits in the core processor as needed. Due to the reduced parallelization potential of the $\pi/8$ operations, we assume a serial scheduling is employed in the QREs presented here. If nonrestrictive availability of magic states is ensured, and considering that the expected time to execute a $\pi/8$ rotation is equal to one logical cycle because of the auto-correcting buffers, the minimum number of logical cycles required to execute the entire circuit in a serial scheduling is equal to its $T$ count, ignoring the warm-up time. Each logical cycle requires performing $d_{\text{core}}$ parity checks, where $d_{\text{core}}$ is the code distance of the logical qubits in the core processor, each taking a time $T_{M}+4T_{2}+2T_{1}+T_{M}+t_{R}$, considering the measurement time $t_{M}$, the reset time $t_{R}$, the single-qubit gate time $t_{1}$, and the two-qubit gate time $t_{2}$. Therefore, while it is easy to generate time estimates for circuits scheduled in serial, generating the space estimates require creating a MSF with enough distillation units capable of distilling magic states quickly enough to keep the core processor constantly busy.
@@
The errors of the core processor and the MSF are modelled and predicted following the pipeline described in Ref. Silva et al. (2024a). Since we provide QREs for the case with a never idling core processor, the accumulated errors in the core are only resulting from the Clifford operations required for the multi-qubit lattice surgeries performed and the protection of the idling data qubits while lattice surgeries involving other data qubits are occurring in the core. The accumulated errors in the core processor is approximated as
$E_{\text{core}}\approx(2Q+\sqrt{8Q}+29)Te_{\text{mem},\text{core}},$ (42)
which assumes that all $(2Q+\sqrt{8Q}+29)$ logical qubits in the core processor (approximated size of the memory fabric and buffer) following the design presented in Figure 15, where $Q$ is the number of data qubits in the circuit, are susceptible to result in an error with probability $e_{\text{mem},\text{core}}$ during all the $T$ logical cycles required to run the circuit. The error rate $e_{\text{mem},\text{core}}$ is derived from emulations of the FTQC protocol for quantum memory. These emulations establish the correlation between code distance and logical error rates based on a given choice of physical parameters, resulting in a predictive model obtained by regression from numerical simulations at low code distances, using efficient stabilizer circuit simulators Gidney (2021) following the descriptions in Section III.2. In the MSF, the error rate of output magic states is resulting from the preparation, distillation and expansion procedures. Following Ref. Silva et al. (2024a), considering $e_{\text{prep}}$ as the error rate for the magic states prepared from physical qubits and that the Clifford and growth accumulated errors can be approximated to the memory errors, i.e., $e_{\text{cliff}}=e_{\text{grow}}=e_{\text{mem}}$, the magic state error rates of the entire MSF using 15:1 distillation units can be calculated recursively as follows:
@@
Therefore, the error rate of the magic state input to the core processor is $e_{\text{core}}=e_{\text{in},L+1}$ for an MSF with $L$ distillation levels. Given that $T$ magic states needs to be distilled for the entire execution of the quantum program, we have
@@
The choice of hardware parameters to determine the required number of distillation levels $L$ and code distances $d_{l},\forall l\in{1,\ldots,L+1}$, which includes the core processor as $l=L+1$, is such that it must reach the target logical error rates based on Equations 41 to 43. The process followed to make these decisions is described in Ref. Silva et al. (2024a). In summary, the core processor’s code distance $d_{L+1}$ is minimized, assuming $E_{\text{msf}}=0$. Then, it sets the first level code distance $d_{1}$ considering the residual error budget left after the core level logical encoding is decided, that is, $E_{\text{msf}}\leq E-E_{\text{core}}$. Next, it determines the number of distillation levels $L$ required to meet the magic state error rate requirement $e_{\text{core}}$ derived from Equation 43 for the residual error budget. Finally, if $L>1$, it calculates the code distances $d_{l}$ for all levels that can meet the error budget using the minimum number of physical qubits across the whole architecture. Once these decisions have been made, the assembler determines the number of distillation units required for a steady flow of magic states to the core processor such that the distillation rate of magic states output from the MSF matches the consumption rate of magic states in the core processor.